### v1

# Synthesizing Strong-Coupling Kohn-Luttinger Superconductivity in 2D Van der Waals materials

Abstract
The Kohn-Luttinger (KL) mechanism of pairing, which describes superconductivity emergent from repulsive interactions, typically yields Cooper pairs at high angular-momentum ($\ell>0$) and extremely low transition temperatures ($T_{c}$). Here, we reveal an inter-layer s-wave ($\ell=0$) KL superconductivity with greatly elevated $T_{c}$ in a multi-layer Hubbard model, which prototypes stacked two-dimensional (2D) electrons in layered van der Waals materials. By employing determinant quantum Monte Carlo and dynamical mean-field theory simulations, we show that a strong pairing attraction $V^{*}$, without the mediation of collective modes, can emerge between inter-layer electrons in the system. As inter-layer repulsion $U$ increases, $V^{*}$ evolves from a conventional KL relation of $V^{*}\propto-U^{2}$, to a linear strong-coupling scaling of $V^{*}\propto-U$, resulting in enhanced superconductivity at large $U$. This strong-coupling KL pairing is robust against changes in lattice geometries and dimensionalities, and it can persist, in the presence of a large remnant Coulomb repulsion $U^{*}$ between pairing electrons. Using ab initio calculations, we propose a few 2D layered van der Waals materials that can potentially realize and control this unconventional superconductivity.
Nickelates superconductivity, Hubbard model, Dynamical mean-field theory, quantum Monte Carlo
preprint: APS/123-QED
Introduction- Van der Waals (VdW) materials hosting two-dimensional electrons have emerged as versatile platforms for engineering and studying many exotic correlated phases of matter, including quantum Hall states [57, 71, 72], correlated Mott insulating states  [4, 73, 54, 24], excitonic insulators [70, 46, 52] , and superconductivity [5, 36, 74, 68, 61, 69, 23, 20, 61]. This field is rapidly advancing due to the unparalleled electrical tunability, broad material palette, and moiré engineering capabilities of VdW materials [49], which continue to promise the discovery of new quantum phases.
On the other hand, in condensed matter physics, the discovering and understanding of new mechanisms of unconventional superconductivity  [3, 62, 28, 76, 42, 15, 33, 48, 11, 56, 8, 12, 31, 14, 64, 75], i.e., in which electron pairing arises sans an electron-phonon-coupling (EPC) origin, have been one of the central research topics. The magnetic-correlation mediated superconductivity (SC) that relevant to cuprates [37, 32, 13, 35], iron-based [9, 66], and nickelates [40, 47, 67, 41] superconductors marks such an example. Another prominent theoretical paradigm of unconventional SC is the Kohn-Luttinger (KL) superconductivity in metals, where overscreened Coulomb interactions can induce Cooper pairs at high angular-momenta ($\ell>0$[27, 26]. Despite its long-standing theoretical proposal, direct experimental evidence for KL SC remains elusive. Recent studies suggest that SC in graphene systems [21, 6, 18], and monolayer $\mathrm{NbSe_{2}}$  [61] could potentially be explained within the KL framework.
In their original proposal, Kohn and Luttinger demonstrated that electron pairing can occur in metals without involving EPC, with $T_{c}/T_{F}\sim exp[-(2\ell)^{4}]$ at large $\ell$, where $T_{F}$ is the Fermi temperature [27]. While such pairing could, in principle, lead to SC, the exponentially suppressed $T_{c}$ makes experimental observation challenging. Subsequent theoretical studies have suggested that lattice effects [2, 53], van Hove singularities [22, 44, 26], the presence of multiple Fermi surfaces [10], and quantum geometry [59, 25] may enhance KL SC. In these studies, KL pairing typically favors a $d$- wave ($\ell=2$[53, 44], or $p$- wave ($\ell=1$) pairing symmetry  [53, 26, 18]. There are also studies suggest that for weakly repulsive electrons in a single-band system, pairing with $\ell<3$ is unlikely to occur [1]. These conclusions are primarily derived from weak-coupling considerations. The behaviour of KL SC at strong-coupling remains unclear to date.
Figure 1: Illustration of the trilayer Hubbard model and KL electron pairing on 2D triangular lattice. (a): AAA-stacking trilayer triangular lattice with spinfull electrons($N_{l}=3$, $N_{\sigma}=2$). Within a unit cell, different Hubbard terms $U_{l\sigma,m\sigma^{{}^{\prime}}}$ are depicted. Long-range interaction $U_{mq}(r)$ terms beyond a unit cell are not shown here. (b): One of the Kohn-Luttinger second order Feynman diagrams leading to inter-layer attraction between a pair of electrons residing on the top- and bottom- layer respectively, via the polorization of the middle-layer. (c): A cartoon showing the inter-layer pairing of two electrons (solid dots) in real-space (indicated by wavy line). Here hollow dots denote holes. For clarity, only one spin species is shown.
In this works, we propose that a trilaminar 2D electron system in stacked van der Waals materials, which can be described by multi-layer Hubbard models (see Fig. 1a), can be utilized to realize an inter-layer $s$- wave strong-coupling KL superconductivity. In the strong-coupling regime, where inter-layer Hubbard repulsion $U$ becomes comparable to electronic bandwidth $W$ ($U\sim W$), the effective pairing attraction $V^{*}$ exhibits a $V^{*}\propto-U$ scaling. This behavior contrasts with the conventional weak-coupling KL pairing where $V^{*}\propto-U^{2}$, and the magnetic-correlation mediated SC in cuprates where $V^{*}\propto-1/U$, $U$ are the respective repulsive coupling constants . Such a linearly growing $V^{*}$ leads to enhanced KL SC at large $U$, which can persists even in the presence of a large Coulomb repulsion $U^{*}$ between paired electrons. We further analyse the dependence of this KL SC on electron densities and long-range Coulomb repulsions. Using density functional theory (DFT) and constrained random phase approximation (cRPA) calculations, we identify several candidate 2D layered half-metals and metals, including metal-doped phosphorene and transition-metal halides [7, 34], as potential platforms for realizing this strong-coupling KL SC.
Model and Methods - We consider a three-layer repulsive Hubbard model as following,
$\mathcal{H}=-t\sum_{\langle i,j\rangle}\sum_{l\sigma}\left(c_{il\sigma}^{\dagger}c_{jl\sigma}^{\phantom{\dagger}}+c_{jl\sigma}^{\dagger}c_{il\sigma}^{\phantom{\dagger}}\right)$
$+\sum_{il\sigma}(\epsilon_{l}-\mu)n_{il\sigma}+\sum_{i}\sum_{l\sigma\neq m\sigma^{\prime}}\frac{U_{l\sigma,m\sigma^{\prime}}}{2}n_{il\sigma}n_{im\sigma^{\prime}}$
$+\sum_{i,r,l,m,\sigma,\sigma^{\prime}}U_{lm}(r)n_{i,l,\sigma}n_{i+r,m,\sigma^{\prime}}$ (1)
where $c^{\dagger}_{il\sigma}$ ($c_{il\sigma}$) denotes the creation (annihilation) operator for a fermion at unit cell $i$, layer $l$, and with spin flavor $\sigma$. The intra-layer nearest-neighbor hopping amplitude $t$ will be used as the energy unit throughout the paper ($t\equiv 1$ ). The electron density $n_{l}=\langle n_{l}\rangle$ in each layer can be tuned by the chemical potential $\mu$ and site-energies $\epsilon_{l}$. Unless specified, we maintain identical density for all three layers: $n_{1}=n_{2}=n_{3}=n$. The Hubbard term $U_{l\sigma,m\sigma^{\prime}}$ describes repulsions between two electrons in the same unit cell with layer and spin indices ($l,\sigma$), and $(m,\sigma^{\prime})$ respectively, $(l,\sigma)\neq(m,\sigma^{\prime})$, $U_{l\sigma,m\sigma^{\prime}}\geq 0$ in our study. $U_{mq}(r)$ denotes intra- or inter-layer long-ranged Coulomb repulsions beyond a unit cell. Here, we focus on two cases on 2D triangular lattice: trilayer spinless model ($N_{l}=3,N_{\sigma}=1$) and trilayer spinfull ($N_{l}=3,N_{\sigma}=2$) model ( See Fig.1). We consider three major term of Hubbard $U_{l\sigma,m\sigma^{\prime}}$: repulsion between electrons at neighboring layers $U\equiv U_{1\sigma,2\sigma^{\prime}}\equiv U_{2\sigma,3\sigma^{\prime}}$, and $U^{*}\equiv U_{1\sigma,3\sigma^{\prime}}$ for electrons on the top- and bottom-layer. There are also on-site repulsions $U_{o}\equiv U_{l\uparrow,l\downarrow}$ in the spinfull model. The effects of non-local $U_{lm}(r)$ will be investigated at the end of the work. To solver Eq. 1, we employ determinant quantum Monte Carlo (DQMC) and dynamical mean-field theory (DMFT) [19] approaches.
Figure 2: Pairing susceptibilities $\chi$ and effective pairing attractions $V^{*}$ of the spinless trilayer Hubbard model on 2D triangular lattice from $6\times 6\times 3$ DQMC simulation. (a): Pairing sueptiblities $\chi_{l}$ for $\ell=s$- wave as a function of temperature $T$ at $U=6,U^{*}=0,n_{1}=n_{2}=n_{3}=n=0.5$. The bubble contributuion $\chi^{0}_{s}$ is shown in dashed line. (b): Effective pairing susceptiblities $\chi_{\mathrm{eff}}=\chi_{l}-\chi_{0,l}$ as a function of temperature $T$. For $p$- and $d$- wave, $\chi_{\mathrm{eff}}<0$ , denoting these pairing symmetries are disfavored by correlation effects. Other parameters are the same as in (a). (c): $\chi_{s}$ and $\chi_{s}^{0}$ as functions of $U^{\prime}$ at $U=6t,n=0.5,T=0.5t$. The critical $U^{\prime}_{c}$ where $\chi_{s}-\chi_{s}^{0}\approx 0$ is used to define the effective pairing attraction $V^{*}=-U^{\prime}_{c}$. (d): $V^{*}$ as a function of $U$ at $n=0.5,T=0.5t$. (e): $V^{*}$ as a function of electron density of the middle layer $n_{2}$ at $U=4,T=0.5$. Here $n_{1}$ and $n_{3}$ are fixed as $n_{1}=n_{3}=0.5$. (f): Inverse $s$- wave pairing susceptibility as a function of temperature $T$. Here $U=6t,n=0.6$. $\chi^{-1}\rightarrow 0$ is indicated as where $T$ approaches superconducting transition temperature $T_{c}$. Here $U^{*}=0$ for all plots.
weak- and strong-coupling limits We first consider tri-layer spinless model with $U_{12}=U_{23}=U,U_{13}=U^{*}=0$, namely, assuming the Coulomb repulsion $U^{*}$ between top- and bottom- layer ($l=3$) are fully screened off. In weak-coupling limit $U\ll t$, the second order KL Feynman diagrams give rise to an inter-layer attraction between top- and bottom- layer (within a unit cell), via the polarization of the middle-layer ($l=2$), see Fig. 1b and Fig. 1c. In the high temperature limit, this effective attraction $V_{\mathrm{eff}}$ scales like (see Appendix),
$V_{\mathrm{eff}}$ $\propto$ $-U^{2}n_{2}(1-n_{2})$ (2)
where $n_{2}$ is the electron filling in the middle-layer. In the large $U/t>>1$ limit at half-filling, to avoid repulsion between electrons (holes) on neighbouring layers, if the top- and bottom- layer are occupied by a pair of electrons (holes), the middle-layer will be occupied by a hole (electron). This occupation configuration has an energy gain being of the order of $U$, hence generating an effective attraction $V_{\mathrm{eff}}$ between top- and bottom- layer electrons,
$V_{\mathrm{eff}}\propto-U.$ (3)
see Fig. 1c and Appendix. Therefore, for both weak-coupling and strong-coupling cases, effective attraction between top- and bottom- layer electrons can be generated from the many-body effects of inter-layer repulsion $U$.
DQMC result - Now we use the numerical exact DQMC to study the spinless model on 2D triangular lattices with $U_{12}=U_{23}=U,U_{13}=U^{*}=0$. In Fig. 2a, we present DQMC result on the pairing susceptibility $\chi=\chi_{\ell}^{(l,m)}$ (see Appendix) for electron pairs on the top- and bottom- layer ($l=1,m=3$) with $\ell$-wave pairing symmetry. In Fig. 2a, $\chi_{s}^{(1,3)}$ is plotted for $\ell=s$ wave as a function of temperature $T$ (solid line) at $U=6,n=0.5$, where one can see that $\chi_{s}^{(1,3)}$ grows rapidly as $T$ decreases. The sign problem prevents DQMC from approaching the BKT temperature $T_{BKT}$ for a diverging pairing susceptiblity $\chi$. A pragmatic approach to tackle the pairing tendency in DQMC is to estimate the effective pairing $\chi_{\mathrm{eff}}=\chi-\chi_{0}$, where $\chi$ and $\chi_{0}$ are respectively the full pairing susceptibility and bubble contribution. The sign of $\chi_{\mathrm{eff}}$ reflects whether the vertex corrections, i.e., the interaction effects beyond single-particle level, enhance superconductivity ($\chi_{\mathrm{eff}}>0$ ) or suppress superconductivity ($\chi_{\mathrm{eff}}<0$ ). Fig. 2a and Fig. 2b suggests that the $s$-wave pairing are favoured by correlation effects as $\chi_{eff}>0$, whereas the $p$- and $d$- wave pair as disfoured since $\chi_{\mathrm{eff}}<0$ for them. We have also considered other pairing correlations, such as the $s$-,p,dp-,d- wave pairings between the top- and middle- layers [$(l,m)=(1,2)$], where they are all disfavored (data not shown).
To quantify the Kohn-Luttinger (KL) pairing mechanism, we introduce an effective attraction strength, $V^{*}$ to describe the attraction between two pairing electrons. This quantity is defined by probing the system’s response to pair-breaking repulsion. Specifically, we calculate $\chi_{\mathrm{eff}}$ for $s$- wave pairing between top- and bottom- layer electrons while introducing an explicit interlayer repulsion $U^{{}^{\prime}}$ between them. For $U^{{}^{\prime}}=0$ , we observe $\chi_{\mathrm{eff}}>0$, indicative of favored pairing, like shown in Fig.2b. Increasing $U^{{}^{\prime}}$ suppresses $\chi_{\mathrm{eff}}$, and we can identify a critical $U_{c}^{{}^{\prime}}$ where $\chi_{\mathrm{eff}}\approx 0$. We then define the effective attraction $V^{*}$ as $V^{*}=-U_{c}^{{}^{\prime}}$, as shown in Fig. 2c. This quantity measures the binding strength between two electrons opposing the pair-breaking repulsion. As shown in Fig. 2d, $V^{*}$ as a function of $U$ follows a quadratic dependence $V^{*}\propto U^{2}$ for $U<5t$, consistent with the conventional KL theory [27, 55].
For larger $U>5t$, the quadratic KL relation no longer holds, $V^{*}$ gradually transits to a new linear scaling, $V^{*}\propto-U$ , underscoring the fundamental difference between KL pairing at strong-couling and the conventional KL SC (dashed line in Fig. 2d). Fig. 2e shows the density dependence of $V^{*}$ at $U=6,T=0.5$, where one sees that the relation $V^{*}\propto n_{e}^{(2)}(1-n_{e}^{(2)})$ is also nicely followed. From above result, we learn that in both weak-coupling and strong-coupling limits, DQMC result of $V^{*}$ agrees well with our analytic analysis on $V_{\mathrm{eff}}$ in Eq. 6 and Eq. 10. In essence, the overall growing tendency of $V^{*}$ with $U$ suggests that one can keep increasing inter-layer Hubbard repulsion $U$ to realize a larger effective attraction $V^{*}$ between electrons, hence potentially a higher transition temperature $T_{c}$ for SC.
Figure 3: Superconducting transition temperature $T_{c}$ from DMFT calculations. (a) $T_{c}$ as a function of $U$ at $n_{1}=n_{2}=n_{3}=0.6$ for a few different $U^{*}$. (b) $T_{c}$ as a function of $U^{*}$ at $U=8t$ at $n=0.6$ (20% electron doping). (C) Same as (b) but at $n=0.7$ (40% electron doping). In (b) and (c) The values of $T_{c}$ at $U^{*}=U/2=4t$ is marked out by dashed lines.
SC instability- We now use CDMFT to compute the superconducting $T_{c}$ at small temperatures. $T_{c}$ is identified as the temperature where pairing susceptibility $\chi$ diverges in DMFT. In Fig. 2f, we plot the inverse of $s$- wave pairing susceptibility, $\chi^{-1}$ as a function of $T$ at $U=6,n=0.6$. DQMC data available at high temperatures are also presented, where excellent agreement between DMFT and DQMC results can be seen. Extrapolating DMFT result of $\chi^{-1}$ to zero, $\chi^{-1}\rightarrow 0$, we obtain a superconducting $T_{c}\approx 0.12t$ for the given parameters in this plot. Repeating the calculations of $T_{c}$ for different $U$, the $U$-dependence of $T_{c}$ is obtained in Fig. 3a. In this plot, one sees that superconducting $T_{c}$ (triangles) grows with $U$ in a broad range of $U$, in consistent with the DQMC result ( $V^{*}$ versus $U$ in Fig. 2d). Specifically, escalated $T_{c}\gtrsim 0.12t$ can be found when $U$ is in the range of $U\sim(6t,10t)$, namely, when $U$ becomes comparable with bandwidth $W$ ($W=9t$ for 2D triangular lattice), KL SC can be greatly enhanced. For comparison, in the magnetic-fluctuation mediated SC in the single-band 2D Hubbard model, the optimal superconducting $T_{c}$ obtained by CDMFT is commonly less than $T_{c}\lesssim 0.05t$[16, 38, 35]. This means that in terms of the reduced pairing energy scale, $T_{c}/t$, the strong-coupling KL SC in the trilayer Hubbard model has an intrinsically stronger pairing force than that in the magnetic-fluctuation mediated SC in Hubbard model.
Figure 4: Phase diagrams showing superconducting instability (SC) and density wave (DW) instability. (a) Superconducting transition temperature $T_{c}$ as a function of electron density $n$ on the 2D triangular lattice at $U=6$. (b) Same as (a) but at $U=12$. (c) Density wave and superconducting transition temperatures as a function of electron density $n$ on the 3D cubic lattice at $U=10$. Here for all three cases $n_{1}=n_{2}=n_{3}=n$. The repulsion between layer- ( spin flavor-) 1 and layer- (spin flavor-) 3, $U_{13}=U^{{}^{\prime}}=0.$
Furthermore, the nonzero $U^{*}$ cases shown in Fig. 3a (dots and diamonds) suggest that in the presence of significant remnant Coulomb repulsion ($U^{*}=U/3,U/2$ ) between paired electrons, large superconducting $T_{c}$ can be still maintained. In Fig. 3b and Fig. 3c, $T_{c}$ as a function of $U^{*}$ are shown at $n=0.6,0.7$ respectively (for fixed $U=8t$), where finite $T_{c}$ can be found even when $U^{*}\geq U/2$. Specifically, for the $n=0.6$ case, an enhanced $T_{c}\approx 0.05t$ is found at $U=8t,U^{*}=4t$. This means that in the large screening length limit $\lambda_{TF}\rightarrow\infty$ along c-axis, i.e., when Coulomb potential in the vertical direction becomes, $U(r_{z})\propto\frac{e^{-r_{z}/\lambda_{TF}}}{r_{z}}\sim\frac{1}{r_{z}}$, (thus $U^{*}\sim U/2$ if layer distances $d_{12}=d_{23}$), the strong-coupling KL SC can still survive. This finding obviously lie beyond the expectation of high-temperature DQMC results presented in Fig. 2d, where $V^{*}\sim-U/2$ for large $U$. In a purely static picture, a condition of $U^{*}\gtrsim U/2$ leads to a net repulsion between top- and bottom-layer electrons [ i.e., $(U^{*}+V^{*})\gtrsim 0$], thereby severely inhibiting pairing between them. Hence, the persistence of a high $T_{c}$ under $U^{*}\gtrsim U/2$ suggests that the retardation part of pairing force may play a role at low temperatures (see discussions in Appendix). A further key point to emphasize is the sizeable $T_{c}$ up to a filling of $n=0.7$ (40% electron doping), as evidenced in Fig. 3c. This presents a clear difference with the case in Hubbard model for cuprates [16], where SC, induced by magnetic correlations, rapidly vanishes beyond $p\lesssim 30\%$ for hole-doping and $p\lesssim 20\%$ for electron-doping. This distinctness is rooted in the “pairing glue”: whereas SC in cuprates relies on the collective spin fluctuations [13] that are suppressed as $n\gg 0.5$ or $n\ll 0.5$ , the strong-coupling KL SC identified here does not involve any intermediate collective modes, thus ensuring robustness across a broad range of $n$.
The phase diagrams - Fig. 4a presents a superconducting phase diagram showing $T_{c}$ as a function of electron density $n$ at $U=6,U^{*}=0$ for the spinless model on 2D triangular lattice from nearly empty ($n\sim 0.2$) to nearly full-filled ($n\sim 0.9$) at $U=6t,U^{*}=0$. One sees that $T_{c}$ is maximumized as $n$ being close to half-filling, in agreement with the $V^{*}$ versus $n$ tendency in DQMC (Fig. 2e). Again, here we find KL SC instability in an extremely large range of $n$. Increasing Hubbard repulsion to $U=12t$ can even further expand the SC regime. For instance, on the electron doped side, KL SC extends to $n_{c}\gtrsim 0.95$ as $T\rightarrow 0$ ( Fig. 4b).
Near half-filling, we find strong density fluctuations in DMFT calculations despite no stable long-range ordered state found on the triangular lattice. For comparison, in Fig. 4c, we present a similar phase diagram for the non-frustrated cubic lattice (where the layer index $l$ now can be understood as a pseudo-spin index). At half-filling ($n=0.5$), this system indeed develops a density wave (DW) phase (Fig. 4c, see also Appendix). This long-range DW instability competes with SC order, which can be destroyed by doping electrons (holes). As DW order vanishes (e.g. $n\gtrsim 0.6$ in Fig. 4c), SC instability emerges. Further increasing $n$ leads to the decreasing of superconducting $T_{c}$, similar to the triangular lattice result. This finding demonstrates that strong-coupling KL SC is robust against changes in lattice geometry (Cubic V.S. Triangular) and dimensionality (3D V.S. 2D).
Figure 5: Effective pairing attraction $V^{*}$ as a function long-range Coulomb repulsion $U_{nn}$ for spinless and spinfull models on 2d Triangular lattice at $U=4,U_{o}=8,U^{\prime}=0,T=1.0$. $U_{nn}$ denotes repulsion between intra-layer nearest-neighbouring sites, $U_{nn}\equiv U_{ll}(|r|=1)$.
Spinful model and effects of $U_{lm}(r)$ - In the spinful model of Eq. 1, twofold competing effects can influence interlayer KL pairing relative to the spinless case: (1) the extra spin degree of freedom in the middle layer can enhance pairing by offering an additional polarization channel mediating attraction; (2) on the other hand, strong on-site repulsion $U_{o}$ between opposite-spin electrons induces local electron–hole pairs ($i.e.$, local moments), which is detrimental to interlayer electron-electron pairing. How $T_{c}$ of KL SC changes in the spinful model is therefore determined by the balance between these two opposing mechanisms.
We examine the influence of both the on-site $U_{o}$ and long-range $U_{lm}(r)$ on $V^{*}$ in Fig. 5. For $U_{lm}(r)$, as a concrete example, we consider its intra-layer nearest-neighbor (NN) component, denoted as $U_{nn}\equiv U_{ll}(|r|=1)$. Fig. 5a shows that at $U_{nn}=0$, the spinfull model (at $U_{o}=8t,U=4t$) yields a significantly smaller effective attraction $V^{*}$ compared to the corresponding spinless model ( dashed line, without $U_{o}$, and $U=4t,U_{nn}=0$), confirming that $U_{o}$ suppresses KL pairing. As $U_{nn}$ increases, $V^{*}$ grows in the spinful case, and it exceeds the spinless value once $U_{nn}>2t$. This finding indicates that $U_{nn}$ in fact boosts KL pairing. This is because long-range Coulomb repulsions generally promotes charge density wave (CDW) fluctuations, which intrinsically suppresses the detrimental local moment effect induced by $U_{o}$. At a higher filling $n=0.8$ (Fig. 5b), $V^{*}$ is also seen enhanced by $U_{nn}$. In this case, the spinful model already gives a larger $V^{*}$ than its spinless counterpart even at $U_{nn}=0$. This is expected, because, far away from half-filling ($n\gg 0.5$), local moments are suppressed by doping, thus the additional polarization channel effect in the spinful model becomes predominant, which results in a large $V^{*}$.
Figure 6: Crystal and band structure of the Iron-doped Phosphorene and decorated transition metal halides $\mathrm{CrCl_{3}}$. (a) Top view of the crystal structure of $\mathrm{P_{35}Fe}$. (b) Energy band structure of $\mathrm{P_{35}Fe}$. (c) Crystal structure of $\mathrm{Na}$-absorbed $\mathrm{CrCl_{3}}$, $\mathrm{NaCr_{2}Cl_{6}}$. (d) Energy band structure of $\mathrm{NaCr_{2}Cl_{6}}$.
Realization of strong-coupling KL pairing- As demonstrated above, enhanced KL pairing can be achieved in a trilayer structure of 2D electrons when interlayer repulsion $U$ is comparable to the bandwidth $W$. In 2D van der Waals layered materials, the interlayer interaction can reach $U\lesssim 0.5eV$ (see Appendix and Ref. [63]). To satisfy the strong-coupling condition $U\gtrsim W$, a narrow electronic bandwidth is therefore needed, $W\lesssim 0.5eV$. Through density functional theory (DFT) and constrained random-phase approximation (cRPA) calculations, we have identified several atomically thin two-dimensional half-metals that can effectively realize the trilayer spinless Hubbard model fulfilling $W\lesssim U$. Examples include iron-doped phosphorene ($\mathrm{FeP_{35}}$), with $W\sim 0.37eV$ and $U\sim 0.15eV$ , and sodium-adsorbed transition-metal trihalides such as $\mathrm{NaCr2Cl_{6}}$, where $W\sim 0.087eV$ and $U\sim 0.3eV$ (see Fig. 6 and Appendix). Finally, recent advances in engineering artificial two-dimensional electron lattices in semiconductor heterostructures have achieved bandwidths of only a few meV [65], offering another promising platform for realizing strong KL superconductivity.
Discussion and conclusion - We have proposed that multi-layered 2D electron systems can be engineered to realize a novel emergent quantum many-body state: strong-coupling KL superconductivity. A key advantage of our proposal lies in the large pairing energy scale of the strong-coupling KL SC, which does not need an intermediate collective mode [64, 75] to mediate the pairing. As a result, it exhibits robustness against variations in lattice geometry and dimensionality, and can endure a large remnant Coulomb repulsion $U^{*}$ between pairing electrons. These advantages allow substantial flexibility to control the pairing by tuning sample thickness, dielectric environment or interlayer distance in VdW materials. Although our low-temperature DMFT calculations of $T_{c}$ are restricted to the spinless model for half-metals due to numerical challenges, the DQMC results do suggest that the spinful model (for normal 2D metals) can, under proper conditions, host an even stronger effective pairing attraction $V^{*}$.
In summary, we have demonstrated the emergence of a robust, strong-coupling inter-layer $s$-wave Kohn-Luttinger superconducting state in multi-layer Hubbard models prototyping stacked 2D electron systems. Our ab initio calculations identify a few candidate materials capable of realizing this novel state, indicating that 2D van der Waals materials can also constitute a versatile platform for designing and manipulating unconventional superconductivity.
Acknowledgment We are grateful to Fadi Sun for useful discussions. This work is supported by the National Natural Science Foundation of China (Grants No.12274472, No. 12494594). We also thank the support from the Research Center for Magnetoelectric Physics of Guangdong Province (Grants No. 2024B0303390001) and Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2401010).