### v2

# Synthesizing Strong-Coupling Kohn-Luttinger Superconductivity in 2D Van der Waals Materials

Abstract
The Kohn-Luttinger (KL) mechanism of pairing, which describes superconductivity emergent from repulsive interactions, typically yields Cooper pairs at high angular-momentum ($\ell>0$) and extremely low transition temperatures ($T_{c}$). Here, we reveal an inter-layer s-wave ($\ell=0$) KL superconductivity with greatly elevated $T_{c}$ in a multi-layer Hubbard model, which prototypes stacked two-dimensional (2D) electrons in layered van der Waals materials. By employing determinant quantum Monte Carlo and dynamical mean-field theory simulations, we show that a strong pairing attraction $V^{*}$, without the mediation of collective modes, can emerge between inter-layer electrons in the system. As inter-layer repulsion $U$ increases, $V^{*}$ evolves from a conventional KL relation of $V^{*}\propto-U^{2}$, to a linear strong-coupling scaling of $V^{*}\propto-U$, resulting in enhanced superconductivity at large $U$. This strong-coupling KL pairing is robust against changes in lattice geometries and dimensionalities, and it can persist, in the presence of a large remnant Coulomb repulsion $U^{*}$ between pairing electrons. Using ab initio calculations, we propose a few 2D layered van der Waals materials that can potentially realize and control this unconventional superconductivity.
Nickelates superconductivity, Hubbard model, Dynamical mean-field theory, quantum Monte Carlo
preprint: APS/123-QED
Introduction- Van der Waals (VdW) materials hosting two-dimensional electrons have emerged as versatile platforms for engineering and studying many exotic correlated phases of matter, including quantum Hall states [57, 73, 74], correlated Mott insulating states  [4, 75, 54, 23], excitonic insulators [71, 46, 52] , and superconductivity [5, 36, 76, 69, 61, 70, 22, 19, 61]. This field is rapidly advancing due to the unparalleled electrical tunability, broad material palette, and moiré engineering capabilities of VdW materials [49, 78], which continue to promise the discovery of new quantum phases.
On the other hand, in condensed matter physics, the discovering and understanding of new mechanisms of unconventional superconductivity  [3, 62, 28, 79, 42, 14, 33, 48, 10, 56, 7, 11, 31, 13, 64, 77], i.e., in which electron pairing arises sans an electron-phonon-coupling (EPC) origin, have been one of the central research topics. The magnetic-correlation mediated superconductivity (SC) that relevant to cuprates [37, 32, 12, 35], iron-based [8, 66], and nickelates [40, 47, 68, 41] superconductors marks such an example. Another prominent theoretical paradigm of unconventional SC is the Kohn-Luttinger (KL) superconductivity in metals, where overscreened Coulomb interactions can induce Cooper pairs at high angular-momenta ($\ell>0$[27, 25]. Despite its long-standing theoretical proposal, direct experimental evidence for KL SC remains elusive. Recent studies suggest that SC in graphene systems [20, 6, 17], and monolayer $\mathrm{NbSe_{2}}$  [61] could potentially be explained within the KL framework.
Figure 1: Illustration of the trilayer Hubbard model and KL electron pairing on 2D triangular lattice. (a): AAA-stacking trilayer triangular lattice with spinfull electrons. Different Hubbard terms $U_{l\sigma,m\sigma^{{}^{\prime}}}$ within a unit cell are depicted. Long-range interaction $U_{mq}(r)$ terms beyond a unit cell are not shown here. (b): A KL second order Feynman diagram that leads to inter-layer attraction between a pair of electrons in the top- and bottom- layer respectively, via the polorization of the middle-layer. (c): A cartoon showing the inter-layer pairing of two electrons /holes (solid/hollow dots) in real-space (indicated by wavy line). The spin label is omitted in (b) and (c) for clarity. (d): Effective pairing attraction $V^{*}$ as a function of $U$ for spinless trilayer Hubbard model from $6\times 6\times 3$ DQMC simulation at $T=0.5,n=0.6$, cf. Fig. 2.
In their original proposal, Kohn and Luttinger demonstrated that electron pairing can occur in metals without involving EPC, below $T_{c}/T_{F}\sim exp[-(2\ell)^{4}]$ at large $\ell$, where $T_{F}$ is the Fermi temperature [27]. While such pairing could, in principle, lead to SC, the exponentially suppressed $T_{c}$ makes experimental observation challenging. Subsequent theoretical studies have suggested that lattice effects [2, 53], van Hove singularities [21, 44, 25], the presence of multiple Fermi surfaces [9], and quantum geometry [59, 24] may enhance KL SC. In these studies, KL pairing typically favors a $d$- wave ($\ell=2$[53, 44], or $p$- wave ($\ell=1$) pairing symmetry  [53, 25, 17]. There are also studies suggest that for weakly repulsive electrons in a single-band system, pairing with $\ell<3$ is unlikely to occur [1]. These conclusions are primarily derived from weak-coupling considerations. The behaviour of KL SC at strong-coupling remains unclear to date.
In this work, we propose that trilaminar 2D electron systems in stacked VdW materials (see Fig. 1a), can be utilized to realize an inter-layer $s$- wave KL superconductivity. In the strong-coupling regime where inter-layer repulsion $U$ becomes comparable to electronic bandwidth $W$ ($U\sim W$), we reveal a $V^{*}\propto-U$ scaling for the effective pairing attraction $V^{*}$. This behavior contrasts with the conventional weak-coupling KL pairing where $V^{*}\propto-U^{2}/t$, and the magnetic-correlation mediated SC in cuprates where $V^{*}\propto-t^{2}/U$, $U$ and $t$ are respectively the electron repulsion and hopping amplitudes. We show that such a linearly growing $V^{*}$ leads to enhanced KL SC at large $U$. We analyse the dependence of pairing on electron densities and long-range Coulomb repulsions. Using density functional theory (DFT) and constrained random phase approximation (cRPA) calculations, we identify several candidate 2D layered half-metals and metals, including transition-metal halides [26], as potential platforms for realizing this strong-coupling KL SC.
Model and Methods - We consider a repulsive Hubbard model on 2D trilayer triangular lattice as following,
$\mathcal{H}=-t\sum_{\langle i,j\rangle}\sum_{l\sigma}\left(c_{il\sigma}^{\dagger}c_{jl\sigma}^{\phantom{\dagger}}+c_{jl\sigma}^{\dagger}c_{il\sigma}^{\phantom{\dagger}}\right)$
$+\sum_{il\sigma}(\epsilon_{l}-\mu)n_{il\sigma}+\sum_{i}\sum_{l\sigma\neq m\sigma^{\prime}}\frac{U_{l\sigma,m\sigma^{\prime}}}{2}n_{il\sigma}n_{im\sigma^{\prime}}$
$+\sum_{i,r,l,m,\sigma,\sigma^{\prime}}U_{lm}(r)n_{i,l,\sigma}n_{i+r,m,\sigma^{\prime}}$ (1)
where $c^{\dagger}_{il\sigma}$ ($c_{il\sigma}$) denotes the creation (annihilation) operator for electrons at unit cell $i$, layer $l$, and with spin flavor $\sigma$. The intra-layer nearest-neighbor hopping amplitude $t$ will be used as the energy unit throughout the paper ($t\equiv 1$ ). The electron density $n_{l}=\langle n_{l}\rangle$ in each layer can be tuned by the chemical potential $\mu$ and site-energies $\epsilon_{l}$. Unless specified, we maintain identical density for all three layers: $n_{1}=n_{2}=n_{3}=n$. The Hubbard term $U_{l\sigma,m\sigma^{\prime}}$ describes repulsions between two electrons in the same unit cell with layer and spin indices ($l,\sigma$) and $(m,\sigma^{\prime})$ respectively, $(l,\sigma)\neq(m,\sigma^{\prime})$, $U_{l\sigma,m\sigma^{\prime}}\geq 0$ in our study. $U_{lm}(r)$ denotes intra- or inter-layer long-ranged Coulomb repulsions beyond a unit cell. Here, we will study both spinless ($N_{\sigma}=1$) and spinful ($N_{\sigma}=2$) cases ( See Fig.1). We consider three major term of Hubbard $U_{l\sigma,m\sigma^{\prime}}$: repulsion between electrons at neighboring layers $U\equiv U_{1\sigma,2\sigma^{\prime}}\equiv U_{2\sigma,3\sigma^{\prime}}$, and $U^{*}\equiv U_{1\sigma,3\sigma^{\prime}}$ for electrons on the top- and bottom-layer. There are also on-site repulsions $U_{o}\equiv U_{l\uparrow,l\downarrow}$ in the spinfull case. The effects of non-local $U_{lm}(r)$ will first be neglected, and be investigated at the end of the work. To solver Eq. 1, we employ determinant quantum Monte Carlo (DQMC) and dynamical mean-field theory (DMFT) [18] approaches.
Figure 2: Pairing susceptibilities $\chi$ and effective pairing attractions $V^{*}$ of the spinless trilayer Hubbard model.(a): Pairing susceptibilities $\chi_{l}$ for $\ell=s$- wave as a function of temperature $T$ at $U=6,U^{*}=0,n=0.5$. The bubble contributuion $\chi^{0}_{s}$ is shown in dashed line. (b): Effective pairing susceptibilities $\chi_{\mathrm{eff}}=\chi_{l}-\chi_{l}^{0}$ as a function of $T$. For $p$- and $d$- wave, $\chi_{\mathrm{eff}}<0$ , denoting both pairing symmetries are unfavoured. Other parameters are the same as in (a). (c): $\chi_{s}$ and $\chi_{s}^{0}$ as functions of $U^{\prime}$ at $U=6t,n=0.5,T=0.5t$. The critical $U^{\prime}_{c}$ where $\chi_{eff}\equiv\chi_{s}-\chi_{s}^{0}\approx 0$ is used to define the effective pairing attraction $V^{*}=-U^{\prime}_{c}$. (d): Inverse $s$- wave pairing susceptibility as a function of temperature $T$. Here $U=6t,n=0.6$. $\chi^{-1}\rightarrow 0$ is indicated as where $T$ approaches superconducting transition temperature $T_{c}$. Here $U^{*}=0$ and $n_{1}=n_{2}=n_{3}=n$ for all plots. Results are obtained by $6\times 6\times 3$ DQMC simulation.
weak- and strong-coupling limits - To gain insight from an analytic perspective, we begin by considering the spinless model. Without loss of generality, we set $U_{12}=U_{23}=U$, and $U_{13}=U^{*}=0,U(r)=0$, i.e., neglecting long-ranged repulsions as well as the direct repulsion between the top and bottom layers. In weak-coupling limit $U\ll t$, second order KL Feynman diagrams yield an inter-layer attraction $V_{\mathrm{eff}}$ between the top- and bottom- layers within the same unit cell, mediated by the polarization of the middle-layer ($l=2$), as illustrated in Fig. 1b. In the instantaneous limit, $V_{\mathrm{eff}}$ scales like (see End Matter),
$V_{\mathrm{eff}}$ $\propto$ $-U^{2}n_{2}(1-n_{2})/t$ (2)
where $n_{2}$ is the electron density of the middle-layer. In the large $U/t>>1$ limit at half-filling, to avoid repulsion between electrons (holes) on neighbouring layers, if the top- and bottom- layer are occupied by a pair of electrons (holes), the middle-layer will be occupied by a hole (electron). This occupation configuration has an energy gain being of the order of $U$, hence generating an effective attraction $V_{\mathrm{eff}}$ between top- and bottom- layer electrons,
$V_{\mathrm{eff}}\propto-U.$ (3)
see Fig. 1c and End Matter. Therefore, in both weak-coupling and strong-coupling regimes, an effective attraction between top- and bottom- layer electrons can be generated from the many-body effects of inter-layer repulsion $U$, which eventually leads to superconductivity in the system.
DQMC result - We now use numerical exact DQMC to study the pairing attraction. We first focus on spinless model with $U_{12}=U_{23}=U,U_{13}=U^{*}=0,U(r)=0$. In Fig. 2a, we present the pairing susceptibility $\chi=\chi_{\ell}^{(l,m)}$ for inter-layer ($l=1,m=3$) $s$-wave pairing (see End Matter) at $U=6,n=0.5$. One can see that as temperature $T$ decreases, $\chi_{s}^{(1,3)}$ (solid line) grows rapidly. The notorious sign problem prevents DQMC from accessing the Berezinskii-Kosterlitz-Thouless (BKT) temperature $T_{BKT}$ for a diverging $\chi$. A pragmatic approach to tackle the pairing tendency in DQMC is to estimate the effective pairing $\chi_{\mathrm{eff}}=\chi-\chi_{0}$, where $\chi$ and $\chi_{0}$ are respectively the full pairing susceptibility and bubble contributions. The sign of $\chi_{\mathrm{eff}}$ reflects whether the vertex corrections, i.e., the interaction effects beyond single-particle level, enhance superconductivity ($\chi_{\mathrm{eff}}>0$ ) or suppress superconductivity ($\chi_{\mathrm{eff}}<0$ ). Fig. 2a and Fig. 2b suggests that the $s$-wave pairing are favoured by correlation effects , whereas the $p$- and $d$- wave are unfavoured. We have also considered other pairing correlations, such as the $s$-,p,dp-,d- wave pairings between the top- and middle- layers [$(l,m)=(1,2)$], where they are all unfavoured (data not shown).
To quantify the attraction between pairing electrons, we compute a quantity $V^{*}$ in DQMC, which is defined by probing the system’s $\chi_{\mathrm{eff}}$ response to pair-breaking repulsion. Specifically, we calculate $\chi_{\mathrm{eff}}$ for $s$- wave pairing between top- and bottom- layer electrons while introducing an explicit interlayer repulsion $U^{{}^{\prime}}$ between them. For $U^{{}^{\prime}}=0$ , one finds $\chi_{\mathrm{eff}}>0$, indicative of favoured pairing, like shown in Fig.2b. Increasing $U^{{}^{\prime}}$ will suppress $\chi_{\mathrm{eff}}$, and a critical $U_{c}^{{}^{\prime}}$ can be identified when $\chi_{\mathrm{eff}}\approx 0$. We then define the effective pairing attraction $V^{*}$ as $V^{*}=-U_{c}^{{}^{\prime}}$, as shown in Fig. 2c. This quantity naturally measures the binding strength between two pairing electrons. As shown in Fig. 1d for the spinless model, our DQMC result of $V^{*}$ exhibits a quadratic $U$ dependence, $V^{*}\propto U^{2}/t$ for small $U$ ($U\lesssim 5t$), in nice agreement with Eq. 2 that complies with the conventional KL theory [27, 55] (see also End Matter).
For larger $U\gtrsim 5t$, the quadratic KL relation no longer strictly holds, $V^{*}$ gradually transits to a new linear scaling, $V^{*}\propto-U$ (dashed line in Fig. 1d), underscoring the fundamental difference between KL pairings at strong-couling and that at weak-coupling. Above results indicate that in both large and small $U$ limits, DQMC result of $V^{*}$ agrees well with our analytic analysis on $V_{\mathrm{eff}}$ in Eq. 2 and Eq. 3. In essence, the overall growing tendency of $V^{*}$ with $U$ suggests that one can keep increasing inter-layer Hubbard repulsion $U$ to realize a larger $V^{*}$ between electrons, hence potentially a higher $T_{c}$ of SC.
Figure 3: Superconducting transition temperature $T_{c}$ from DMFT calculations. (a) $T_{c}$ as a function of $U$ at $n_{1}=n_{2}=n_{3}=0.6$ for a few different $U^{*}$. (b) $T_{c}$ as a function of $U^{*}$ at $U=8t$ at $n=0.6$ (20% electron doping). (C) Same as (b) but at $n=0.7$ (40% electron doping). In (b) and (c) The values of $T_{c}$ at $U^{*}=U/2=4t$ is marked out by dashed lines.
SC instability- We now use CDMFT to compute the superconducting $T_{c}$ at low temperatures. $T_{c}$ is identified as the temperature where pairing susceptibility $\chi$ diverges in CDMFT. In Fig. 2d, we plot the inverse of $s$- wave pairing susceptibility, $\chi^{-1}$ as a function of $T$ at $U=6,n=0.6$. DQMC data available at high temperatures are also presented, where excellent agreement between DMFT and DQMC results can be seen. Extrapolating CDMFT result of $\chi^{-1}$ to zero, $\chi^{-1}\rightarrow 0$, we obtain a superconducting $T_{c}\approx 0.12t$ for the given parameters in this plot. Repeating the calculations of $T_{c}$ for different $U$, the $U$-dependence of $T_{c}$ is obtained in Fig. 3a. In this plot, one sees that $T_{c}$ (triangles) grows with $U$ in a broad parameter range, in consistent with the growing tendency of $V^{*}$ in DQMC result ( Fig. 1d). Specifically, escalated $T_{c}\gtrsim 0.12t$ can be found when $U$ is in the range of $U\sim(6t,10t)$, namely, when $U$ becomes comparable with bandwidth $W$ ($W=9t$ for 2D triangular lattice), KL SC can be greatly enhanced. For comparison, in the magnetic-fluctuation mediated SC in the single-band 2D Hubbard model, the optimal superconducting $T_{c}$ obtained by CDMFT is commonly less than $T_{c}\lesssim 0.05t$[15, 38, 35]. This means that in terms of the reduced pairing energy scale, $T_{c}/t$, the strong-coupling KL SC in the trilayer Hubbard model has an intrinsically stronger pairing force than that in the magnetic-fluctuation mediated SC in Hubbard model.
Figure 4: Phase diagrams showing superconducting instability (SC) and density wave (DW) instability. (a) Superconducting transition temperature $T_{c}$ as a function of electron density $n$ on the 2D triangular lattice at $U=6t$. (b) Same as (a) but at $U=12t$. (c) Density wave and superconducting transition temperatures as a function of electron density $n$ on the 3D cubic lattice at $U=10t$. Here for all three cases $n_{1}=n_{2}=n_{3}=n$. The repulsion between layer- (pseudo-spin ) 1 and layer- (pseudo-spin) 3, $U^{*}=0.$
Furthermore, the nonzero $U^{*}$ cases shown in Fig. 3a (dots and diamonds) suggest that in the presence of significant remnant Coulomb repulsion ($U^{*}=U/3,U/2$ ) between paired electrons, large superconducting $T_{c}$ can be still maintained. In Fig. 3b and Fig. 3c, $T_{c}$ as a function of $U^{*}$ are shown at $n=0.6,0.7$ respectively (for fixed $U=8t$), where finite $T_{c}$ can be found even when $U^{*}\geq U/2$. Specifically, for the $n=0.6$ case, an enhanced $T_{c}\approx 0.05t$ is found at $U=8t,U^{*}=4t$. This means that in the large screening length limit $\lambda_{TF}\rightarrow\infty$ along $c$-axis, i.e., when Coulomb potential in the vertical direction becomes $U(r_{z})\propto\frac{e^{-r_{z}/\lambda_{TF}}}{r_{z}}\sim\frac{1}{r_{z}}$, (thus $U^{*}\sim U/2$ if layer distances $d_{12}=d_{23}$), the strong-coupling KL SC can still survive. This finding lie beyond the expectation of high-temperature DQMC results presented in Fig. 1d, where $V^{*}\sim-U/2$ for large $U$. In a purely static picture, $U^{*}\gtrsim U/2$ will lead to a net repulsion between top- and bottom-layer electrons [ i.e., $(|V^{*}|-U^{*})\lesssim 0$], thereby severely inhibiting pairing between them. We attribute the persistence of high $T_{c}$ under $U^{*}\gtrsim U/2$ condition to the retarded pairing or suppression of layer density fluctuations (see discussions in End Matter). A further remarkable point to emphasize is the sizeable $T_{c}$ up to a filling factor of $n=0.7$ (40% electron doping), as shown in Fig. 3c. This presents a clear difference with cases in Hubbard model for cuprates [15], where SC rapidly vanishes beyond $p\lesssim 30\%$ for hole-doping and $p\lesssim 20\%$ for electron-doping. This distinctness is rooted in the “pairing glue”: SC in cuprates relies on the collective spin fluctuations [12], which are suppressed at large dopings. In contrast, the strong-coupling KL SC identified here does not involve any intermediate collective modes, thus ensuring robustness across a broad range of $n$.
The phase diagrams - Fig. 4a presents a superconducting phase diagram showing $T_{c}$ as a function of electron density $n$ for the spinless model on 2D triangular lattice from nearly empty ($n\sim 0.2$) to nearly full-filled ($n\sim 0.9$) at $U=6t,U^{*}=0$. One sees that $T_{c}$ is maximumized as $n$ being close to half-filling. Again, here we find SC instability in an extremely large range of $n$. Increasing repulsion $U$ can even further expand the SC regime. For instance, on the electron doped side, finite $T_{c}$ can be found for $n_{c}\gtrsim 0.95$ at $U=12t$ ( Fig. 4b).
In vicinity of half-filling, we find strong density fluctuation in CDMFT calculations which tends to form staggered distribution of electrons at the three layers ( see Fig. 5b ). However, due to the geometry frustration, no true long-range ordered state is found on triangular lattice. For comparison, in Fig. 4c, we present a similar phase diagram for the non-frustrated cubic lattice (where the layer index $l$ now can be understood as a pseudo-spin index). At half-filling ($n=0.5$), this system indeed develops a density wave (DW) phase. This long-range DW instability competes with SC order, which can be destroyed by doping electrons (holes). As DW order vanishes (e.g. $n\gtrsim 0.6$ in Fig. 4c), SC instability emerges. Further increasing $n$ leads to the decreasing of superconducting $T_{c}$, similar to the triangular lattice result. This finding demonstrates that strong-coupling KL SC is robust against changes in lattice geometry (bipartite V.S. non-bipartite) and dimensionality (3D V.S. 2D).
Figure 5: Long-range Coulomb repulsions can enhance pairing attraction $V^{*}$. $U_{nn}=U_{ll}(\vec{r_{1}})$ denotes the intra-layer NN repulsion, $\vec{r_{1}}||_{xy}=1$.(a) $V^{*}$ as a function of $U_{nn}$ for spinless model at $T=1.0$. Other NN repulsions are set as $U_{12}(\vec{r_{1}})=U_{23}(\vec{r_{1}})=U_{nn}/2$, $U_{13}(\vec{r_{1}})=U_{nn}/4$. (b) A staggered distribution of electrons at three layers minimizing the potential energy is shown in 1D chain, on which interaction terms that promote (solid lines) or suppress (dashed lines) inter-layer pairing is illustrated. Filled (hollow) symbols denote electrons (holes). (c) and (d): $V^{*}$ as a function of $U_{nn}$ for the spinful model at $U_{o}=8,U=4,U^{*}=0,T=1.0$. Inter-layer NN repulsion $U_{lm}(\vec{r_{1}}),l\neq m$ is neglected in (c) and (d) to avoid sign problem.
Fffects of $U(r)$ and spinful model- Considering long-ranged Coulomb repulsions beyond a unit cell, there can be different $U_{lm}(r)$ terms stabilizing or destabilizing the binding of electron (hole) pairs at top- and bottom- layers, as illustrated in Fig. 5b. We find that the overall effects of the nearest-neighboring (NN) repulsion enhances $V^{*}$, as shown in Fig. 5a for the spinless model.
For the spinful model of Eq. 1, twofold competing effects can influence interlayer KL pairing relative to the spinless case: (1) the extra spin degree of freedom in the middle layer can enhance pairing by offering an additional polarization channel; (2) on the other hand, strong on-site repulsion $U_{o}$ between opposite-spin electrons induces local electron–hole pairs ($i.e.$, local moments), which is detrimental to interlayer electron-electron pairing. The balance between these two opposing mechanisms then affects $T_{c}$ in the spinful model.
We examine influence of both on-site $U_{o}$ and NN repulsion $U_{nn}$ on $V^{*}$ in Fig. 5c and Fig. 5d at $U_{o}=8t,U=4t$. Fig. 5c shows that at $U_{nn}=0$, the spinfull model yields a smaller $V^{*}$ compared to the corresponding spinless model ( dashed line, without $U_{o}$, and $U=4t,U_{nn}=0$), confirming that $U_{o}$ suppresses KL pairing. As $U_{nn}$ increases, $V^{*}$ grows and exceeds the spinless value once $U_{nn}\gtrsim 2t$. We note that long-range Coulomb repulsions can promote charge density wave (CDW) fluctuations, hence suppressing the detrimental local moments induced by $U_{o}$. This could be another route that $U(r)$ enhances $V^{*}$ in spinful cases. At a higher filling $n=0.8$ (Fig. 5b), $V^{*}$ is also seen enhanced by $U_{nn}$. In this case, the spinful model already gives a larger $V^{*}$ than its spinless counterpart even at $U_{nn}=0$. This is expected, because, far away from half-filling, local moments are suppressed by doping, thus the additional polarization channel effect in spinful model becomes predominant, which results in a larger $V^{*}$.
Figure 6: Crystal and band structure of $\mathrm{Na}$-adsorbed transition metal halides $\mathrm{CrCl_{3}}$. (a) Crystal structure of $\mathrm{Na}$-adsorbed $\mathrm{NaCr_{2}Cl_{6}}$. (b) Energy band structure of $\mathrm{NaCr_{2}Cl_{6}}$. Inset: Magnified view of energy bands near Fermi level.
Realization of strong-coupling KL pairing- As demonstrated above, enhanced KL pairing can be achieved in a trilayer structure of 2D electrons when $U\sim W$. In 2D van der Waals layered materials, the interlayer repulsion can reach $U\lesssim 0.5eV$ (see End Matter and Ref. [63]). To satisfy the strong-coupling condition, a narrow electronic bandwidth, $W\lesssim 0.5eV$, is therefore needed. Through density functional theory (DFT) and constrained random-phase approximation (cRPA) calculations, we have identified several atomically thin two-dimensional half-metals that can effectively realize the trilayer spinless Hubbard model fulfilling $W\lesssim U$. Examples include sodium-adsorbed transition-metal trihalides such as $\mathrm{NaCr_{2}{C}l_{6}}$, where $W\sim 0.087eV$ and $U\sim 0.3eV$ (see Fig. 6 and End Matter), and an iron-doped phosphorene, $\mathrm{FeP_{35}}$ [34] (see End Matter). We note that the transition metal dichalcogenide(TMD)-based moiré superlattices [78] exhibit bandwidths below $10\mathrm{meV}$ (inter-site hopping strength $t\sim 1\mathrm{meV}$[67], which is smaller than the inter-layer repulsion [72]. Hence these systems may also be suitable for implementing our proposal. Finally, recent advances in engineering artificial 2D electron lattices within semiconductor heterostructures have yielded bandwidths of only a few meV [65], offering another promising platform for realizing strong KL superconductivity.
Discussion and conclusion - We have proposed a scheme for realizing a novel emergent quantum many-body state: strong-coupling KL superconductivity in 2D electrons. Recently, novel exciton-mediated SC in atomically thin VdW heterostructures have been proposed [64, 77]. A key distinction of our proposal lies in the large pairing energy scale $V^{*}$ that does not require any collective mode to mediate the pairing. Consequently, it exhibits robustness against variations in lattice geometry and dimensionality, and can endure a large remnant Coulomb repulsion $U^{*}$ between pairing electrons. These advantages allow substantial flexibility to control SC by tuning sample thickness, dielectric environment or interlayer distance in VdW materials. Although our low-temperature DMFT calculations of $T_{c}$ are restricted to the spinless model for half-metals due to numerical challenges, the DQMC results do suggest that the spinful model (for normal 2D metals) can, under proper conditions, host an even stronger effective pairing attraction $V^{*}$.
In summary, we have demonstrated the emergence of a robust, strong-coupling inter-layer $s$-wave Kohn-Luttinger superconducting state in multi-layer Hubbard models prototyping stacked 2D electron systems. Our ab initio calculations identify a few candidate materials capable of realizing this novel state, indicating that 2D van der Waals materials can also constitute a versatile platform for designing and manipulating unconventional superconductivity.
Acknowledgment We are grateful to Wenchen Luo and Fadi Sun for useful discussions. This work is supported by the National Natural Science Foundation of China (Grants No.12274472, No. 12494594). We also thank the support from the Research Center for Magnetoelectric Physics of Guangdong Province (Grants No. 2024B0303390001) and Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2401010). H.Y. acknowledges support by Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2501003) and NSFC under Grant No. 12274477.